OpenStudy (anonymous):

When 1618 J of heat energy is added to 38.8 g of ethanol, C2H6O, the temperature increases by 17.0 °C. Calculate the molar heat capacity of C2H6O.

OpenStudy (abb0t):

start by calculating the molar mass for EtOH \(\sf \color{red}{NOTE}\)\(\sf \color{}{:~EtOH = ethanol}\) Once you get the molar mass, use dimensional analysis to cancel out units, grams, g: \(\sf \color{}{g~of~EtOH \times \frac{mol~of~EtOH}{mass~EtOH}}\), and then you'll use the formula for molar heat capacity, which I believe is: \(\sf \color{blue}{C_p=\frac{q_p}{n \Delta T}}\), where q\(_p\) = heat energy, n = mole (which you found above), and \(\Delta\)T = change in temp. And there you have it. Molar heat capacity for EtOH (ethanol)

OpenStudy (anonymous):

Thank you!